5 g of a gas is contained in a rigid container and is heated from 15°C to 25°C. Specific heat capacity of the gas at constant volume is 0.172 cal/g-°C and the mechanical equivalent of heat is 4.2 J/cal. Calculate the change in the internal energy of the gas.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. m = 5 g, ∆t = 25 – 15 = 10°C CV = 0.172 cal/g-°CJ = 4.2 J/Cal. dQ = du + dw Now, V = 0 (for a rigid body) So, dw = 0. So dQ = du. Q = msdt = 5 × 0.172 × 10 = 8.6 cal = 8.6 × 4.2 = 36.12 Joule.